OpenStudy (anonymous):

Barb ate 6 cheeseburgers in the last 4 months. how many do u predict she will eat in two years?

sam (.sam.):

6/4=3/2 cheeseburgers per month In 2 years, 3/2 * 24month=36 cheeseburgers

OpenStudy (anonymous):

thank you